the perimeter of a field is 140 feet. the width is one-sixth of the length. find the width of the field.
Answer by phxladyphx@yahoo.com(18)

p=140 feet
x=width
6x=length
x+x+6x+6x=140
14x=140
14x/14=140/14
x=10
The width is 10 feet and the length is 6x so 6(10) which is 60 feet.